# Exploring Neural Granger Causality with xLSTMs: Unveiling Temporal Dependencies in Complex Data

### v4

@@ Appendix D Extended Experimental Results
This section supplements the experimental findings of Section 4.1. Specifically, Figure 7 provides the full set of relations extracted from the Company Fundamentals dataset.
Figure 7: GC-xLSTM uncovers relationships between economic variables of individual companies.
Appendix E Additional Insights into Training
Figure 8(a) shows how the different loss components change during training. The robustness of training is reflected in the stability of variable usage once an optimal set is found, as Figure 8(b) shows.
@@ Appendix E Additional Insights into Training
Figure 8: Algorithm 1 jointly optimizes the prediction loss while adaptively establishing sparsity. Results show training on Lorenz-96 with $F=40$ and $T=1000$.
Appendix F Scaling Behaviour
Figure 9 shows the training time and peak GPU memory reserved during the training of GC-xLSTM on the Lorenz-96 dataset with $T=1000$ for various numbers of variates $V$.

### v5

@@ Appendix D Extended Experimental Results
This section supplements the experimental findings of Section˜4.1. Specifically, Figure˜7 provides the complete set of relations extracted from the Company Fundamentals dataset. On this challenging full feature set, only some of the discovered GC relations are economically plausible.
Figure 7: GC-xLSTM uncovers relationships between economic variables within individual companies. Established GC links are highlighted in dark blue.
Appendix E Scaling Behaviour
Figure˜8 shows the training time and peak GPU memory reserved during the training of GC-xLSTM on the Lorenz-96 dataset with $T=1000$ for various numbers of variates $V$.
Appendix F Additional Insights into Training
Figure˜9(a) shows how the different loss components change during training. The robustness of training is reflected in the stability of variable usage once an optimal set is found, as Figure˜9(b) shows.
@@ Appendix F Additional Insights into Training
Figure 9: Algorithm˜1 jointly optimizes the prediction loss while adaptively establishing sparsity. Results show training on Lorenz-96 with $F=40$ and $T=1000$.